OpenStudy (anonymous):

factor the trinomial 25r^2+5r-42

OpenStudy (anonymous):

so you have to think of mltiplies of 25 and 42 that would give me 5 25 is 25*1 or 5*5 42is 42*1, 2*21, 3*14,6*7 we look at 5*6=30 and 7*5= 35 then 35-30= 5 ok (__x __ __) (__x __ __)=(5x__6)(5x__7) so what goes in the blanks minus or plus